OpenStudy (anonymous):

You roll two dice What is the probability that the sum of the dice is odd or the sum of the dice is greater than 6?

OpenStudy (anonymous):

sum of dice is greater than 6={(1,6),(2,5),(2,6),(3,4),(3,5),(3,6)...(6,6)}=21 P(A)=21/36=7/12 sum is odd={(1,2),(1,4),(1,6),(2,1),(2,3),(2,5),(3,2),(3,4),...(6,5)}=18 P(B)=18/36=1/2 P(A intersection B)={(1,6),(2,5),(3,4),(3,6),(4,3),(4,5),(5,2),(5,4),(5,6),(6,1),(6,3),(6,5)}=12 12/36=1/3 P(AUB)=7/12+1/2-1/3 hope its correct